# Cosmology with voids

### v3

@@ 3.2 The Zel’dovich approximation
A related analytical approach for tracking void growth is the Zel’dovich approximation (Zel’dovich, 1970; Hidding et al., 2013; White, 2014), which connects linear and nonlinear phases of structure formation. The Zel’dovich approximation quantifies the displacement of a point particle at some initial position by the peculiar gravity of density fluctuations in the approximation that these fluctuations grow according to linear perturbation theory. For example, a particle sitting at some location in the primordial density field is accelerated by the gravity of fluctuations around it. Its subsequent motion is in the direction of that acceleration. At early times when linear theory applies, only the magnitude of the acceleration changes. Thus the motion of the particle is simply along a single displacement vector. The approximation is formally dependent on linear theory, but in a Lagrangian or ballistic framework that empowers it beyond linear theory (White, 2014).
@@
$\vec{\psi}(\vec{x})\sim\vec{\nabla}\phi;$ (24)
here, $\phi$ is the peculiar gravitational potential, related to the density contrast $\delta(\vec{x})$ (equation (10)) through the Poisson equation. In the Fourier domain, this expression is simpler (White, 2014):
$\vec{\psi}=\frac{1}{(2\pi)^{3}}\int d^{3}k\frac{i\vec{k}}{k^{2}}\delta(\vec{k})\exp^{i\vec{q}\cdot\vec{k}},$ (25)
where $\delta(\vec{k})$ refers to Fourier modes of the density contrast. In the approximation, these modes all grow at the same rate (equation (12)), and the scaling of their collective amplitudes determines the time when the displacement is evaluated. For example, when $\delta$ is scaled to the present epoch, as for the primordial power spectrum in figure 1, $\vec{x}$ gives the location of the particle at the present epoch.
@@
$\dot{\vec{x}}=Hf\vec{\psi},$ (26)
where $H$ is the Hubble parameter and $f$ is the velocity growth factor from linear theory (equation (18)).
@@
Because of its extraordinary effectiveness beyond the linear regime, the “quasilinear” Zel’dovich approximation works well for voids (Yoshisato et al., 2006), a conclusion that has been compellingly validated for voids as small as 10 $h^{-1}$pc with numerical simulation (e.g., Schuster et al., 2023).
@@ 3.3 Nonlinear theory: The shell model
To determine the evolution of voids with depths that are of order unity, formally beyond the reach of linear perturbation theory and even the Zel’dovich approximation, we perform nonlinear calculations that more accurately track the flow of mass. Following the strategy of Blumenthal et al. (1992), we model a void as a series of thin, concentric shells, each with an assigned mass uniformly distributed within it. Newtonian mechanics provides the framework for solving the detailed evolution of these shells over distances that are much smaller than the Hubble length, $c/H_{0}\approx 4,000$ Mpc.
In our implementation, we adopt initial conditions as in Blumenthal et al. (1992), with a starting redshift of $z_{i}\gg 1$, and a spherical, constant-density region that is underdense relative to the rest of the universe by some small factor, $\varepsilon_{\text{init}}$, at that redshift. The initial mass density is
$\rho_{i}(x)=\bar{\rho}\left[1+\delta(x)\right]\ \hskip 115.63243pt$
$=\frac{3H_{0}\Omega_{\text{M}}}{8\pi Ga_{i}^{3}}\times\begin{cases}1-\varepsilon_{\text{init}}&\text{if}\ x<x_{v}\\ 1&\text{otherwise}\end{cases}$ (27)
@@
For the initial velocities, we assume that matter closely tracks the peculiar flow of growing density perturbations as predicted by linear theory (Peebles, 1980). We set the initial proper radial velocity of each shell according to equation (17), which gives the outflow speed as a function of distance from the void center. Since shells are very thin compared with the void radius, we just use the average of the shell’s inner and outer radius for the distance.
@@
which follows from a Newtonian representation of cosmic acceleration, taking full advantage of spherical symmetry (Peebles, 1980). As in Blumenthal et al. (1992), we choose mass elements as thin, concentric shells centered on the underdense region. Each shell contains mass in accordance with its initial volume and local mass density (equation (3.3)). The radial velocity comes from equation (17). We track the trajectories of these shells by solving the equation of motion (equation (28)) numerically with Python’s solve_ivp routine in the SciPy.integrate module.
@@
There are other approaches to tracking void evolution in the idealized case, taking full advantage of the spherical symmetry and the assumption of uniform density within a void (Sheth & van de Weygaert, 2004; Pace et al., 2010; Wagner et al., 2015; Dai et al., 2015; Desjacques et al., 2018). The shell model is straightforward to implement and has flexibility to manage more complicated density profiles, modified force laws, and any shell crossings.
Figure 2 illustrates of the paths of shells in an extremely deep void forming within a $\Lambda$CDM model. The example tracks shells in an underdense region with an initial outer comoving radius $x_{v}=100$ Mpc and constant $\epsilon=0.1$ at $z_{i}=1000$, with no peculiar motion at that starting redshift. We choose these parameters to facilitate comparison with figure 1 in Blumenthal et al. (1992), showing calculations with similar initial conditions. Although this underdense region is unrealistically deep for voids of any scale, much less for one of this large size, its evolution poses a clear, challenging nonlinear problem as a numerical benchmark.
Figure 2: Mass flow lines in a uniform low-density region as a function of the scale factor $a$ for a $\Lambda$CDM cosmological model with starting conditions similar to Blumenthal et al. (1992). The density inside the region is 10% less than that of the surrounding material at $z=1000$ and all mass elements initially follow the Hubble flow. Each gray line tracks the path of a thin spherical shell. The shells inside the low-density region expand more rapidly than those outside and quickly overtake them. The black curve shows the boundary between the emerging void and the background model where a high-density zone builds and persists.
@@ 3.4 Comparing approaches to idealized void evolution
$\epsilon(r)=\left(1-\frac{2r^{2}}{3r_{v}}\right)\exp^{-r^{2}/r_{v}^{2}},$ (29)
@@ 3.5 The evolution of idealized voids
Figure 5 gives another view of the impact of both the mass density and dark energy in terms of void depth, $\varepsilon$, in equation‘(15), ranging from 0 (no void) to 1 (vacuum). The models shown in curves run from $\Omega_{\text{M}}=0$ to $\Omega_{\text{M}}=1$, both with and without a cosmological constant ($\Omega_{\Lambda}=1-\Omega_{\text{M}}$ and $\Omega_{\Lambda}=0$, respectively). In contrast to earlier plots, this figure tracks voids with sizes and depths comparable to those observed in the cosmic web.
@@
Figure 6: The outflow rate inside a void versus void depth for several cosmological models and at two redshifts. The legend lists the shell simulations; the dark gray curves show the analytical approximations from equation (30) for the outflow rate, formally defined as the Hubble parameter inside a void in units of the background Hubble parameter. Light gray curves are predictions from linear theory.
@@
$\frac{{H_{\text{eff}}}}{H}=1+\frac{f\varepsilon\left(1+0.22\varepsilon\right)}{3},$ (30)
@@
where the void depth $\varepsilon$ is the measured value. This expression, derived from fitting to a subset of cosmological models, extends equation (22) to voids with depths up to $\sim 0.5$, accurate to within 0.3% of the nonlinear results in all the cosmological models explored here. Coupled with accurate approximations to the growth parameters (Lahav et al., 1991),
$D\approx\frac{{\Omega_{\text{M}}(z)}}{(1+z)\Omega_{\text{M}}}\ \times$
$\frac{\Omega_{\text{M}}^{4/7}-\Omega_{\Lambda}+(1+\Omega_{\text{M}}/2)(1+\Omega_{\Lambda}/70)}{{\Omega_{\text{M}}(z)}^{4/7}-{\Omega_{\Lambda}(z)}+[1+{\Omega_{\text{M}}(z)}/2][1+{\Omega_{\Lambda}(z)}/70]},$ (31)
$f\approx\left[\frac{(1+z)^{3}\Omega_{\text{M}}H_{0}^{2}}{H(z)^{2}}\right]^{0.6},$ (32)
@@
the approximation in equation (30) leads to computationally fast predictions of void properties from cosmological parameters. This feature is important to enable the dense sampling required by model parameter fits that we use extensively below.
@@
We also present an analytical approximation for the void depth $\varepsilon$, determined from nonlinear calculations, as a function of the linear theory prediction, $\varepsilon_{\text{lin}}$ (equation (16)):
$\varepsilon=\varepsilon_{\text{lin}}\left[1-0.60\varepsilon_{\text{lin}}\left(1-0.27\varepsilon_{\text{lin}}\right)\right].$ (33)
@@
This expression, like equation (30), was also derived from fitting with a subset of cosmological models. It is accurate to within 3% for $\varepsilon\lesssim 0.5$ in all the cosmological models considered here. (See reference (Bernardeau, 1994) for an elegant derivation of a similar relationship.)
As with void outflow rate and depth, void radius is not accurately predicted by linear theory. In linear theory, the void radius is constant in comoving coordinates as material within it streams out into the surrounding universe. The prediction is that voids deepen but do not expand relative to the Hubble flow. In contrast, a true void may lose mass from its interior as in linear theory, but it also grows in size relative to the comoving background. If a void has no mass loss from its interior, then it evolves only through geometric expansion. Its radius is predicted from the mass it contains, $\sim 1-\varepsilon$, and that fact that this mass is conserved. For a given void depth, this radius is an upper bound. If some mass flows out of the void, the void does not grow as much to achieve that same depth. We obtain a lower bound to the void size by assuming that the mass lost from the void’s interior is as high as the linear theory prediction. Together, these bounds provide this constraint on void radius:
$\left(\frac{\varepsilon_{\text{lin}}}{\varepsilon}\right)^{1/3}\lesssim\frac{R_{\text{co}}}{R_{\text{co}}(z_{\text{init}})}\lesssim\left(\frac{1}{1-\varepsilon}\right)^{1/3}$ (34)
where $R_{\text{co}}$ is the actual (nonlinear) void radius in comoving coordinates, $R_{\text{co}}(z_{\text{init}}$ is the initial comoving radius, $\varepsilon$ is the actual void depth, and $\varepsilon_{\text{lin}}$ is the depth from linear theory, derived from the initial underdensity augmented by the density growth factor $D$. Our numerical experiments with the shell model confirm that void radii lie between the two bounds.
@@ 3.6 The largest voids
To determine the linearity scale $\lambda_{\text{lin}}$, we measure the statistic $\sigma_{\text{R}}$, the typical relative density fluctuation amplitude in a tophat window function of radius $R$, given the power spectrum of a cosmological model (equation (13)). The models we consider are normalized so that the linearity scale is somewhat larger than 8 $h^{-1}$pc at a redshift $z=0$; for “1-$\sigma$” fluctuations in a $\Lambda$CDM universe, $\sigma_{\text{R}}$ dips below 0.2 at a length of just under 40 $h^{-1}$pc. However, voids this large may arise from more rare, deeper fluctuations. In a volume of radius $R_{\text{3G}}\equiv 3,000$ $h^{-1}$pc, there are approximately $R_{\text{3G}}^{3}/4R^{3}\sim 10^{4}$ non-overlapping samples of the density field convolved with a tophat of radius $R=40$ $h^{-1}$pc. Assuming Gaussian statistics and statistical independence of samples, we expect a “one-of-a-kind” fluctuation of $3.7\times 0.2$, or, equivalently, a void depth of over 0.7, well beyond the reach of linear theory (figure 3). Thus $\lambda_{\text{lin}}$ must be considerably larger.
@@ 3.7 Cues from idealized voids
Simulations of idealized cosmic voids based on the shell model and equation (28) provide a testing ground for understanding voids and their connection to the background cosmology. The main results are:
@@
  • Void depth and outflow rate are simply connected by an analytical expression, equation (30), a modest extension to linear theory. This expression allows estimates of $fH$ (the velocity growth parameter times the Hubble parameter of the background universe) as a function of redshift, independent of void radius.
The analytical expression in equation (30) is a central result. In principle, with samples of $fH$ over a range of redshifts, we can estimate cosmological parameters in a manner similar to studies of Type 1a supernova (Kirshner, 1999). We next amplify this result.
@@ 4 Cosmological parameter estimation from IDEALIZED voids
The interior regions of idealized voids evolve as isolated mini-universes. They have FLRW parameters of their own. Nonetheless there are two key connections between voids and the background cosmology. The first connection is the void depth $\varepsilon$, which relates the matter density inside a void to the mean density of the universe. The second connection is equation (30), which links $\varepsilon$ and the outflow rate ${H_{\text{eff}}}$ to $fH$, where $f$ is the velocity growth parameter (equation (18)) and $H$ is the background Hubble parameter. Just as samples of $H(z)$ from Type Ia supernovae yield $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$, samples of $fH$ from void properties $\varepsilon$ and ${H_{\text{eff}}}$ can provide cosmological constraints.
Here we infer the parameters of the background cosmology from properties of idealized voids. Our approach is based on sets of numerically evolved idealized voids with depths and outflow rates measured at multiple redshifts. We compare these measurements and the predictions of cosmological models from the approximation for void growth in equation (30). To quantify this comparison, we adopt a likelihood function, based on the sum of square residuals between measured void properties and model predictions. We use a maximum likelihood algorithm to find best-fit parameters of the background cosmology. Because the approximation in equation (30) is not exact, we include a low level of artificial noise in the analysis so that this small difference does not affect our interpretations. Maps of the likelihood function in parameter space reveal the general sensitivity of parameter estimation to sampling strategies that include the choice of the number of voids and the redshifts when they are sampled. This approach allows a demonstration that simple void properties, when measured for just a few voids over a range of redshifts, provide robust constraints on the background cosmology.

### v4

@@ 3.2 The Zel’dovich approximation
A related analytical approach for tracking void growth is the Zel’dovich approximation [91, 107, 92], which connects linear and nonlinear phases of structure formation. The Zel’dovich approximation quantifies the displacement of a point particle at some initial position by the peculiar gravity of density fluctuations in the approximation that these fluctuations grow according to linear perturbation theory. For example, a particle sitting at some location in the primordial density field is accelerated by the gravity of fluctuations around it. Its subsequent motion is in the direction of that acceleration. At early times when linear theory applies, only the magnitude of the acceleration changes. Thus the motion of the particle is simply along a single displacement vector. The approximation is formally dependent on linear theory, but in a Lagrangian or ballistic framework that empowers it beyond linear theory [92].
@@
$\vec{\psi}(\vec{x})\sim\vec{\nabla}\phi;$ (3.10)
here, $\phi$ is the peculiar gravitational potential, related to the density contrast $\delta(\vec{x})$ (equation (2.10)) through the Poisson equation. In the Fourier domain, this expression is simpler [92]:
$\vec{\psi}=\frac{1}{(2\pi)^{3}}\int d^{3}k\frac{i\vec{k}}{k^{2}}\delta(\vec{k})\exp^{i\vec{q}\cdot\vec{k}},$ (3.11)
where $\delta(\vec{k})$ refers to Fourier modes of the density contrast. In the approximation, these modes all grow at the same rate (equation (2.12)), and the scaling of their collective amplitudes determines the time when the displacement is evaluated. For example, when $\delta$ is scaled to the present epoch, as for the primordial power spectrum in figure 1, $\vec{x}$ gives the location of the particle at the present epoch.
@@
$\dot{\vec{x}}=Hf\vec{\psi},$ (3.12)
where $H$ is the Hubble parameter and $f$ is the velocity growth factor from linear theory (equation (3.4)).
@@
Because of its extraordinary effectiveness beyond the linear regime, the “quasilinear” Zel’dovich approximation works well for voids [108], a conclusion that has been compellingly validated for voids as small as 10 $h^{-1}$pc with numerical simulation, e.g., [70].
@@ 3.3 Nonlinear theory: The shell model
To determine the evolution of voids with depths that are of order unity, formally beyond the reach of linear perturbation theory and even the Zel’dovich approximation, we perform nonlinear calculations that more accurately track the flow of mass. Following the strategy in reference [19], we model a void as a series of thin, concentric shells, each with an assigned mass uniformly distributed within it. Newtonian mechanics provides the framework for solving the detailed evolution of these shells over distances that are much smaller than the Hubble length, $c/H_{0}\approx 4,000$ Mpc.
In our implementation, we adopt initial conditions as in [19], with a starting redshift of $z_{i}\gg 1$, and a spherical, constant-density region that is underdense relative to the rest of the universe by some small factor, $\varepsilon_{\text{init}}$, at that redshift. The initial mass density is
$\rho_{i}(x)=\bar{\rho}\left[1+\delta(x)\right]=\frac{3H_{0}\Omega_{\text{M}}}{8\pi Ga_{i}^{3}}\times\begin{cases}1-\varepsilon_{\text{init}}&\text{if}\ x<x_{v}\\ 1&\text{otherwise}\end{cases}$ (3.13)
@@
For the initial velocities, we assume that matter closely tracks the peculiar flow of growing density perturbations as predicted by linear theory [84]. We set the initial proper radial velocity of each shell according to equation (3.3), which gives the outflow speed as a function of distance from the void center. Since shells are very thin compared with the void radius, we just use the average of the shell’s inner and outer radius for the distance.
@@
which follows from a Newtonian representation of cosmic acceleration, taking full advantage of spherical symmetry [84]. As in [19], we choose mass elements as thin, concentric shells centered on the underdense region. Each shell contains mass in accordance with its initial volume and local mass density (equation (3.13)). The radial velocity comes from equation (3.3). We track the trajectories of these shells by solving the equation of motion (equation (3.14)) numerically with Python’s solve_ivp routine in the SciPy.integrate module.
@@
There are other approaches to tracking void evolution in the idealized case, taking full advantage of the spherical symmetry and the assumption of uniform density within a void [109, 110, 111, 112, 113]. The shell model is straightforward to implement and has flexibility to manage more complicated density profiles, modified force laws, and any shell crossings.
Figure 2 illustrates of the paths of shells in an extremely deep void forming within a $\Lambda$CDM model. The example tracks shells in an underdense region starting at a redshift of $z_{i}=1000$ with an outer comoving radius $x_{v}=100$ Mpc, a constant $\epsilon=0.1$, and peculiar velocity here set to $\dot{x}=0$ everywhere. Although the evolution of a spherical tophat void in this Newtonian approximation is scalable to any radius, we choose these specific starting conditions to facilitate comparison with figure 1 in reference [19], showing calculations with similar initial conditions. Although this underdense region is unrealistically deep for voids of any scale, much less for one of this large size, its evolution poses a clear, challenging nonlinear problem as a numerical benchmark.
Figure 2: Mass flow lines in a uniform low-density region as a function of the scale factor $a$ for a $\Lambda$CDM cosmological model with starting conditions similar to reference [19]. The density inside the region is 10% less than that of the surrounding material at $z=1000$ and all mass elements initially follow the Hubble flow. Each gray line tracks the path of a thin spherical shell. The shells inside the low-density region expand more rapidly than those outside and quickly overtake them. The black curve shows the boundary between the emerging void and the background model where a high-density zone builds and persists.
@@ 3.4 Comparing approaches to idealized void evolution
$\epsilon(r)=\left(1-\frac{2r^{2}}{3r_{v}}\right)\exp^{-r^{2}/r_{v}^{2}},$ (3.15)
@@ 3.5 The evolution of idealized voids
Figure 5 gives another view of the impact of both the mass density and dark energy in terms of void depth, $\varepsilon$, in equation‘(3.1), ranging from 0 (no void) to 1 (vacuum). The models shown in curves run from $\Omega_{\text{M}}=0$ to $\Omega_{\text{M}}=1$, both with and without a cosmological constant ($\Omega_{\Lambda}=1-\Omega_{\text{M}}$ and $\Omega_{\Lambda}=0$, respectively). In contrast to earlier plots, this figure tracks voids with sizes and depths comparable to those observed in the cosmic web.
@@
Figure 6: The outflow rate inside a void versus void depth for several cosmological models and at two redshifts. The legend lists the shell simulations; the dark gray curves show the analytical approximations from equation (3.16) for the outflow rate, formally defined as the Hubble parameter inside a void in units of the background Hubble parameter. Light gray curves are predictions from linear theory.
@@
$\frac{{H_{\text{eff}}}}{H}=1+\frac{f\varepsilon\left(1+0.22\varepsilon\right)}{3},$ (3.16)
@@
where the void depth $\varepsilon$ is the measured value. This expression, derived from fitting to a subset of cosmological models in table 2, extends equation (3.8) to voids with depths up to $\sim 0.5$, accurate to within 0.3% of the nonlinear results in all the cosmological models listed in the table. Coupled with accurate approximations to the growth parameters [114],
$D\approx\frac{{\Omega_{\text{M}}(z)}}{(1+z)\Omega_{\text{M}}}\ \cdot\ \frac{\Omega_{\text{M}}^{4/7}-\Omega_{\Lambda}+(1+\Omega_{\text{M}}/2)(1+\Omega_{\Lambda}/70)}{{\Omega_{\text{M}}(z)}^{4/7}-{\Omega_{\Lambda}(z)}+[1+{\Omega_{\text{M}}(z)}/2][1+{\Omega_{\Lambda}(z)}/70]},$ (3.17)
$f\approx\left[\frac{(1+z)^{3}\Omega_{\text{M}}H_{0}^{2}}{H(z)^{2}}\right]^{0.6},$ (3.18)
@@
the approximation in equation (3.16) leads to computationally fast predictions of void properties from cosmological parameters. This feature is important to enable the dense sampling required by model parameter fits that we use extensively below.
@@
We also present an analytical approximation for the void depth $\varepsilon$, determined from nonlinear calculations, as a function of the linear theory prediction, $\varepsilon_{\text{lin}}$ (equation (3.2)):
$\varepsilon=\varepsilon_{\text{lin}}\left[1-0.60\varepsilon_{\text{lin}}\left(1-0.27\varepsilon_{\text{lin}}\right)\right].$ (3.19)
@@
This expression, like equation (3.16), was also derived from fitting with a subset of cosmological models. It is accurate to within 3% for $\varepsilon\lesssim 0.5$ in all the cosmological models considered here. (See reference [115] for an elegant derivation of a similar relationship.)
As with void outflow rate and depth, linear theory does not accurately predict void radius. In linear theory, a void deepens as mass within it streams through its outer boundary, but the void’s comoving radius, $R_{\text{co}}$, does not change; its size is tied to the Hubble flow. In another view, a void acts as an isolated mini-universe. The void’s precursor, with depth $\varepsilon_{\text{init}}$ and comoving radius $R_{\text{init}}$ at some high redshift $z_{\text{init}}$, expands more rapidly than the Hubble flow, increasing in depth as a result. A more realistic scenario involves both mass loss and radial expansion. To include both effects, we use the precursor’s depth to predict $\varepsilon_{\text{lin}}$, the void depth from linear theory (equation (2.11)) to find an upper bound to the mass lost from the void. Assuming that the difference between $\varepsilon_{\text{lin}}$ and the true void depth $\varepsilon$ stems from the additional expansion of the void, we get a lower bound to $R_{\text{co}}$. Combining these results,
$\left(\frac{\varepsilon_{\text{lin}}}{\varepsilon}\right)^{1/3}\lesssim\frac{R_{\text{co}}}{R_{\text{init}}}\lesssim\left(\frac{1}{1-\varepsilon}\right)^{1/3},$ (3.20)
where the lower limit is our estimate with mass loss from the void interior, while the upper limit is from the isolated void model and mass conservation. Our numerical experiments with the shell model confirm that void radii lie between these two bounds.
@@ 3.6 The largest voids
To determine the linearity scale $\lambda_{\text{lin}}$, we measure the statistic $\sigma_{\text{R}}$, the typical relative density fluctuation amplitude in a tophat window function of radius $R$, given the power spectrum of a cosmological model (equation (2.13)). The models we consider are normalized so that the linearity scale is somewhat larger than 8 $h^{-1}$pc at a redshift $z=0$; for “1-$\sigma$” fluctuations in a $\Lambda$CDM universe, $\sigma_{\text{R}}$ dips below 0.2 at a length of just under 40 $h^{-1}$pc. However, voids this large may arise from more rare, deeper fluctuations. In a volume of radius $R_{\text{3G}}\equiv 3,000$ $h^{-1}$pc, there are approximately $R_{\text{3G}}^{3}/4R^{3}\sim 10^{4}$ non-overlapping samples of the density field convolved with a tophat of radius $R=40$ $h^{-1}$pc. Assuming Gaussian statistics and statistical independence of samples, we expect a “one-of-a-kind” fluctuation of $3.7\times 0.2$, or, equivalently, a void depth of over 0.7, well beyond the reach of linear theory (figure 3). Thus $\lambda_{\text{lin}}$ must be considerably larger.
@@ 3.7 Cues from idealized voids
Simulations of idealized cosmic voids based on the shell model and equation (3.14) provide a testing ground for understanding voids and their connection to the background cosmology. The main results are:
@@
  • Void depth and outflow rate are simply connected by an analytical expression, equation (3.16), a modest extension to linear theory. This expression allows estimates of $fH$ (the velocity growth parameter times the Hubble parameter of the background universe) as a function of redshift, independent of void radius.
The analytical expression in equation (3.16) is a central result. In principle, with samples of $fH$ over a range of redshifts, we can estimate cosmological parameters in a manner similar to studies of Type 1a supernova [116]. We next amplify this result.
@@ 4 Cosmological parameter estimation from idealized voids
The interior regions of idealized voids evolve as isolated mini-universes. They have FLRW parameters of their own. Nonetheless there are two key connections between voids and the background cosmology. The first connection is the void depth $\varepsilon$, which relates the matter density inside a void to the mean density of the universe. The second connection is equation (3.16), which links $\varepsilon$ and the outflow rate ${H_{\text{eff}}}$ to $fH$, where $f$ is the velocity growth parameter (equation (3.4)) and $H$ is the background Hubble parameter. Just as samples of $H(z)$ from Type Ia supernovae yield $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$, samples of $fH$ from void properties $\varepsilon$ and ${H_{\text{eff}}}$ can provide cosmological constraints.
Here we infer the parameters of the background cosmology from properties of idealized voids. Our approach is based on sets of numerically evolved idealized voids with depths and outflow rates measured at multiple redshifts. We compare these measurements and the predictions of cosmological models from the approximation for void growth in equation (3.16). To quantify this comparison, we adopt a likelihood function, based on the sum of square residuals between measured void properties and model predictions. We use a maximum likelihood algorithm to find best-fit parameters of the background cosmology. Because the approximation in equation (3.16) is not exact, we include a low level of artificial noise in the analysis so that this small difference does not affect our interpretations. Maps of the likelihood function in parameter space reveal the general sensitivity of parameter estimation to sampling strategies that include the choice of the number of voids and the redshifts when they are sampled. This approach allows a demonstration that simple void properties, when measured for just a few voids over a range of redshifts, provide robust constraints on the background cosmology.